OpenStudy (anonymous):

1. Are 24, 45, and 51 the side measurements of a special right triangle? Yes or No, Explain your answer and show work 1. Are 24, 45, and 51 the side measurements of a special right triangle? Yes or No, Explain your answer and show work @Mathematics

OpenStudy (anonymous):

Yes, 24^2*45^2 = 51^2 576 + 2025 = 2601

OpenStudy (anonymous):

can u explain how u did tht ?

OpenStudy (anonymous):

It is the Pythagoras theorem. Surely when we multiply 2 small numbers we would get the large numbers. So we take 24 and 45. According to pythagoras theorem a^2+b^2=c^2. So, 24^2+45^2=51^2 Check whether it satisfies the above equation. If it is perfectly satisfied then those are measurements of a Right Triangle.

OpenStudy (anonymous):

ooooooooooooh ok i get it thank u

OpenStudy (anonymous):

\[Hypotenuse ^{2} = Sum of \square of legs of that \right \triangle\]

OpenStudy (anonymous):

ok well u got 2601 how do we know its a special right triangle?